Which of the ordered pairs
(3, 1), (0, –4), (–4, 0), (–3, –7)
are solutions for the equation x – y = 4?
I came up with (–4, 0), (–3, –7)as my answers. I think I am right.
-4-0=4
-3-(-7)=4

Answer by Edwin McCravy(20086)


Which of the ordered pairs 
(3, 1), (0, –4), (–4, 0), (–3, –7)
are solutions for the equation x – y = 4?

Let's plug them all in and see:

------------------------
Plug in (x, y) = (3, 1)
 
x – y = 4
      ?
3 - 1 = 4
      ?
    2 = 4

Nopers, (3, 1) is not a solution to the
equation x - y = 4
----------------------------------

Plug in (x, y) = (0, -4)
 
   x – y = 4
         ?
0 - (-4) = 4
         ?
       4 = 4

Yeppers, (0, -4) is a solution to the
equation x - y = 4
----------------------------------
Plug in (x, y) = (-4, 0)
 
   x – y = 4
         ?
(-4) - 0 = 4
         ?       
      -4 = 4

Nopers, (-4, 0) is not a solution to the
equation x - y = 4
----------------------------------
Plug in (x, y) = (–3, –7)
 
      x – y = 4
            ?
(-3) - (-7) = 4
            ?
     -3 + 7 = 4
            ?
          4 = 4
Yeppers, (-3, -7) is a solution to the
equation x - y = 4
----------------------------------

I came up with (–4, 0), (–3, –7)as my answers. 
I think I am right.
-4-0=4
-3-(-7)=4

Yeppers, I think you are right, too!

Edwin